Could someone help me with this one please. If you do not mind would you show me step by step how you came upon the answer? Thank you
Solve (x-5)^2=6 Answer by stanbon(75887) 
 Solve (x-5)^2=6Solve (x-5)^2=6
----------
Take the square root of both sides to get:
x-5 = sqrt6 or x-5 = -sqrt6
Add 5 to both sides to get:
x = 5+sqrt6 or x= 5-sqrt6
=============
Cheers,
stan H.